Fermi National Laboratory


Accelerator Update for September 9, 2005 - September 12, 2005

Friday, September 9
The day shift began with Operations monitoring the Recycler stash and the Antiproton Source (Pbar) stack, with the Tevatron (TeV) in shot setup, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store, and with MiniBooNE and NuMI taking beam.

Operations established combination store 4378 at 10:04 AM with an initial luminosity of 89.64E30. Pbar resumed stacking and the Meson experiment MCenter resumed taking beam.

NuMI experts began a one-hour target scan at 10:21 AM.

Operations put the Recycler into shot setup at 2:44 PM to stash antiprotons.

The Recycler stash ended and Pbar resumed stacking at 3:24 PM. MCenter also resumed taking beam.

A Main Injector (MI) expert began a damper study at 3:36 PM.

The Recycler began preparing to stash antiprotons at 7:39 PM. The stash ended at 8:45 PM and Pbar resumed stacking and MCenter resumed taking beam.

Saturday, September 10
The midnight shift began with Operations monitoring store 4378 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing to stash antiprotons at 12:56 AM. The stash ended at 1:50 AM. Pbar resumed stacking and MCenter resumed taking beam.

At 6:19 AM, Operations reported losing store 4378, the stack, the stash, and the Meson beam permits; Linac also went down and many Booster RF stations tripped off. The cryo system technicians at CHL reported that their lights flickered just before the Tevatron (TeV) quenched at sector B0.

Operations re-established beam to MiniBooNE at 6:57 AM.

A High Voltage expert reported at 7:27 AM that a ComEd transformer and buss in Itasca, IL faulted and caused our power glitch.

Operations established beam to NuMI at 7:31 AM.

At 7:40 AM, MCenter experimenters reported that they couldn’t reset their analyzing magnets (MC7AN1 & 2).

At 8:26 AM, power supply experts began work on an Accumulator quadrupole power supply (A:LQ).

Operators prepared CDF for access at 9:40 AM. They came out of access about an hour and a half later.

At 9:46 AM, MI suffered a vacuum burst in their 52 region due to a septa tripping off during a move. Operators disabled MCenter beam while they investigated the septa problems

D0 went into an hour access at 10:01 AM.

Operators had the MI septa back on by 10:20 AM. MCenter resumed taking beam at 10:23 AM.

At 10:26 AM, at RF expert began working on a Debuncher RF station (DRF1-3). He completed the work in thirty minutes.

At 10:54 AM, D:QF tripped off. A power supply expert investigated.

Operations reported that Pbar resumed stacking at low intensity at 11:10 AM. Normal stacking resumed at 11:30 AM.

A Recycler expert began a study period at 12:26 PM.

Operations put the TeV through a dry squeeze at 1:20 PM.

At 2:24 PM, D:QF tripped off for the first of two times in about an hour. It reset each time.

Operations put the TeV through a wet squeeze at 2:28 PM. A TeV expert began a study with this injection of protons.

The Recycler expert completed his study at 5:51 PM.

At 7:21 PM, the Pbar beam permit for the AP1 power supplies tripped off. It reset eventually.

TeV studies ended at 7:40 PM.

Operation put the TeV into shot setup around 7:48 PM.

Sunday, September 11
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

Operations established store 4384 at 12:48 AM with an initial luminosity of 62.2E30. Pbar resumed stacking and MCenter resumed taking beam.

At 3:23 AM D:QF tripped off for the first of five times in about four hours. It reset each time.

The Recycler began preparations to stash antiprotons at 5:24 AM. The stash ended and Pbar resumed stacking at 6:40 AM. MCenter also resumed taking beam.

The Recycler began stashing more antiprotons at 11:04 AM. The stash ended and stacking resumed and MCenter resumed taking beam at 11:54 AM.

At 12:07 PM, Operations reported that D:QF failed. This halted stacking. Operators called in power supply experts. The experts had the supply repaired by 3:09 PM, but Operations was unable to get Pbar to resume stacking. Beam would not circulate in the Debuncher. Operations called in Pbar experts.

At 3:46 PM, NuMI tripped off due to a problem with a power supply (V118). Operators and a power supply expert managed to bring the supply back on at 4:23 PM. There was a problem with its current-meter.

At 5:11 PM, E:V118 tripped off again. A power supply expert investigated and had the supply up and running by 5:48 PM. He reported that it was an inside breaker problem.

Operations prepared the Pbar tunnel for access at 7:56 PM to investigate. They found damage to the flags and bolts on D1Q13. Operations called in techs to make magnet repairs. Below is a picture of one of the bolts. The other bolt was found on the floor in pieces.

Pbar experts reported that the magnet and RF repairs were completed by 10:47 PM.

Pbar resumed stacking at 11:15 PM.

Monday, September 12
The midnight shift began with Operations monitoring store 4384 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began stashing at 1:34 AM and ended at 2:20 AM. Pbar and MCenter resumed normal operations.

Operations terminated store 4384 at 3:27 AM.

Operation put the TeV into shot setup at 3:40 AM.

Operations reported at 4:10 AM that DRF1-7 had tripped off during the Pbar access. An expert came in later in the morning and fixed the station.

Operations established store 4385 at 6:09 AM with an initial luminosity of 65.6E30. Pbar and MCenter resumed normal operations.

The Plans for Monday
The plans for today are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/12/2005   email Fermilab